# On the Parallelizability of Approval-Based Committee Rules

### v1

@@ Abstract
Approval-Based Committee (ABC) rules are an important tool for choosing a fair set of candidates when given the preferences of a collection of voters. Though finding a winning committee for many ABC rules is NP-hard, natural variations for these rules with polynomial-time algorithms exist. The Method of Equal Shares, an important ABC rule with desirable properties, is also computable in polynomial time. However, when working with very large elections, polynomial time is not enough and parallelization may be necessary. We show that computing a winning committee using these ABC rules (including the Method of Equal Shares) is P-hard, thus showing they cannot be parallelized. In contrast, we show that finding a winning committee can be parallelized when the votes are single-peaked or single-crossing for the important ABC rule Chamberlin-Courant.
@@ 1 Introduction and Related Work
Finding a winning committee for many ABC rules is NP-hard (LeGrand et al., 2007; Procaccia et al., 2008; Aziz et al., 2015; Skowron et al., 2016; Godziszewski et al., 2021; Brill et al., 2024). However, natural variations of these NP-hard rules exist where a winning set of candidates can be found in polynomial time. In addition, the Method of Equal Shares, which has the desirable property of electing a committee that satisfies Extended Justified Representation, can be computed in polynomial time (Peters and Skowron, 2020).
@@
It is important that we can always easily compute the outcome of an election. Previous work on studying the complexity of winner determination has largely focused on whether the winner problem is in P or NP-hard, equating being in P with being easy to compute. However, when elections are very large it is not enough to just have a polynomial-time algorithm. In this case parallelization may be necessary.
@@
The parallelizability of voting rules was previously considered for single-winner voting rules (Csar et al., 2018, 2017), but not for the case of multiwinner rules. In fact, the recent textbook on ABC rules by Lackner and Skowron (2023) explicitly includes as open question “Q20” to determine which polynomial-time ABC rules are inherently sequential. As pointed out by Lackner and Skowron (2023), Approval Voting (AV) is clearly parallelizable (and the analogous result for Satisfaction Approval Voting (SAV) follows from the same argument). We show that computing the winning committee for all other studied polynomial-time ABC rules is inherently sequential.
@@
To show that a problem is inherently sequential we use the notion of P-hardness, since it is generally assumed that P-hard problems are not parallelizable (Greenlaw et al., 1995). We mention that this approach was used previously in computational social choice to show results about single-winner elections (specifically that winner determination for resolute versions of STV and Ranked Pairs is P-complete) (Csar et al., 2017, 2018), problems related to the aggregation of CP-nets (Lukasiewicz and Malizia, 2022), and for finding the essential set (Brandt and Fischer, 2008).
@@
When studying the complexity of a voting problem it is natural to consider settings where the votes of the electorate have structure. Two important domain restrictions are single-peaked and single-crossing preferences (Black, 1948; Mirrlees, 1971), which each model settings where the votes of the electorate are based on a single dimension (e.g., liberal vs. conservative). Both of these models have natural counterparts for approval voting (Faliszewski et al., 2011; Elkind and Lackner, 2015). We show that for single-peaked and for single-crossing approval votes finding a winning committee using the important Chamberlin-Courant ABC rule can be parallelized. We note here that the only previously-known parallelizability results for voting rules that we are aware of (in addition to AV and SAV mentioned previously (Lackner and Skowron, 2023)) are the one for Schultze Voting (Csar et al., 2018) as well as computing the Copeland, Smith, and Schwartz choice sets (Brandt et al., 2009).
@@
Inherently sequential ABC rules. We provide P-hardness reductions to prove the following theorem which encompasses all studied polynomial-time ABC rules from (Lackner and Skowron, 2023) except AV and SAV which are easily seen to be parallelizable (Lackner and Skowron, 2023); in fact they are computable in logarithmic space (see Section A of the appendix for completeness).
@@
Computing a winning committee is inherently sequential for seq-CC, seq-PAV, rev-seq-CC, rev-seq-PAV, seq-Phragmén, Greedy Monroe, and the Method of Equal Shares.
@@
Instead of providing parallel algorithms, we show membership in the relatively unknown complexity class OptL from Álvarez and Jenner (1993) (Definition  14). This allows for stronger results (since OptL is thought to be a strict subset of the class of parallelizable problems) and simpler proofs (since describing parallel algorithms can be an arduous task, e.g., one has to decide on a model and describe how each machine aggregates information, etc.).
@@ 2.1 Approval-Based Committee (ABC) Rules
Given candidates $C$, collection of votes $V$, and committee size $k$, output a committee $W$ of size $k$ such that the most votes have at least one approved candidate in the committee.
@@ 2.2 Computational Complexity
Most of our results concern showing problems to be inherently sequential. It is widely believed that problems that are P-hard are inherently sequential (Greenlaw et al., 1995). Note that when showing a problem is hard for a given complexity class we must be careful to use a reduction with less computational power than the class we are showing hardness for. As is standard we use logspace reductions for our P-hardness results, meaning that the reduction is computable in logarithmic space. [^1] Though logarithmic space may seem quite restrictive, standard NP-hardness reductions are computable in logarithmic space.
@@
It is easy to see that FNC is a subset of FP, the function counterpart of P. Our results that show problems inherently sequential are based on the generally accepted assumption that ${\rm FNC}\neq{\rm FP}$, and we show inclusion in FNC for parallelizability.
@@ 3 Inherently Sequential ABC Rules
Recall that finding the winning committee is a function problem. However, it gives more insight to prove the hardness of a related decision problem. Thus, we consider the decision problem that asks about the containment of a single candidate. First, we show that this formulation is correct by showing the equivalence of the decision and function problem with respect to parallelizability.
@@ 3.1 Chamberlin-Courant
While computing the winning committee for CC is NP-hard (Procaccia et al., 2008) in the general case, simple greedy approaches have been introduced to mimic CC in polynomial time. One such approach is the sequential Chamberlin-Courant rule, denoted seq-CC, where we start with an empty committee and iteratively add the candidate that increases the number of currently hit votes the most (see (Lackner and Skowron, 2023)).
@@
We will show that computing the winning committee using seq-CC is P-hard. Specifically, we show that deciding whether a given candidate is picked for the seq-CC winning committee is P-hard. To do this, we give a logspace reduction from a closely related P-complete problem, OVR, defined below. We note that as with NP-hardness reductions, half the battle is finding the right problem to reduce from, and we were able to find a problem almost equivalent to seq-CC. We hope that the the straightforward nature of our reduction will ease the reader into P-hardness reductions.
@@
Figure 1: Column (a) shows the graph input to OVR at the top and the corresponding seq-CC election at the bottom. Columns (b) and (c) show the state of OVR and seq-CC after one and two rounds, respectively. Candidates are underlined when they are put in the winning committee and votes are darkened when they are hit. Observe how the removal of the highest-degree vertex in OVR corresponds to its inclusion in the winning committee in seq-CC.
@@ 3.2 Method of Equal Shares
The Method of Equal Shares (MES) is a recently introduced rule (Peters and Skowron, 2020) that allots voters budgets and allows them to elect candidates using shares of their budget. This rule has many desirable properties. In particular, it satisfies Extended Justified Representation (a notion of group fairness) and is computable in polynomial time. A generalized version of this rule (Peters et al., 2021), has also been used in the real world for participatory budgeting (a setting that generalizes multiwinner elections to voting over issues to be funded) (see (Peters and Skowron, 2024)).
@@
$x_{r+1}(i)=\begin{cases}x_{r}(i)-\rho_{c}&\text{if $i$ approves $c$ and $x_{r}(i)\geq\rho_{c}$}\\ 0&\text{if $i$ approves $c$ and $x_{r}(i)<\rho_{c}$}\\ x_{r}(i)&\text{otherwise}\end{cases}$
Note that voters may run out of the budget necessary to elect another candidate before $k$ rounds. In this case, the method terminates early and returns the winning committee constructed so far. However, technically this does not meet our definition of an ABC rule, since we must return a committee of size $k$. To rectify this, another ABC rule is called to select the remaining candidates. Lackner and Skowron (2023) suggest using seq-Phragmén (see Sections C.2 and C.4 of the appendix) since it preserves some desirable properties.
We will show that computing the committee found by MES is P-hard. This shows that the hardness of MES is not caused by the hardness of the rule called afterward (such as seq-Phragmén). As in the previous section, we focus on the decision problem where we ask about the inclusion of a specified candidate. We reduce from a problem known as LFMIS:
Problem 10 (Lexicographically First Maximal Independent Set (LFMIS) (Cook, 1985)).
@@
Computing the winning committee for MES is inherently sequential.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidate $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by MES includes $c$ if and only if $v$ is included in the LFMIS of $G$.
@@
We note that the reduction above may return a winning committee with fewer than $k$ candidates. Recall that in this case another ABC rule is called to elect the remaining candidates. For the sake of completeness, we show in Section C.4 of our appendix that MES is also P-hard when seq-Phragmén is used, as suggested by Lackner and Skowron (2023), to finish electing a total of $k$ candidates. Particularly, we reduce from LFMIS and show that the specified vertex $v$ is picked by MES if and only if $v$ belongs to the LFMIS.
@@ 4 Domain Restrictions
Figure 3: Column (a) shows an election. Column (b) shows how the votes are single-peaked with respect to the axis $c\ b\ a\ d$. Column (c) shows how the votes are single-crossing with respect to the axis 1 2 3 412341\ 2\ 3\ 41 2 3 4. The intervals (candidates) in Column (c) are ordered in nondecreasing order with respect to starting point, as will be done in the construction of Theorem 16.
@@
As pointed out by Faliszewski et al. (2011) (for single-peaked) and Elkind and Lackner (2015) (for single-crossing), the problem of determining whether a collection of votes (in our setting of approval ballots) is single-peaked or single-crossing, and computing an axis witnessing this is in essence the problem of determining whether a 0-1 matrix has the consecutive ones property and computing a permutation that witnesses that. These problems have long been known to be computable in polynomial time (Fulkerson and Gross, 1965; Booth and Lueker, 1976).
@@
We provide a visualization in Figure 4. It is known that every OptL function is parallelizable (Álvarez and Jenner, 1993, Theorem 4.2). We will show that SP-CC and SC-CC are in OptL (Theorems 15 and 16), which immediately implies Theorem 13.
@@
For $1\leq i\leq n$, let $[s_{i},t_{i}]$ be the interval corresponding to the $i$th voter, i.e., $s_{i}$ and $t_{i}$ are in $\{1,\ldots,m\}$ and the $i$th voter approves the candidates from the $s_{i}$th candidate to the $t_{i}$th candidate on our SP axis $L$. Note that we can’t store all these values, but we can recompute each value in logarithmic space whenever it is needed.
We now need to compute a set of $k$ candidates that hit (have nonempty intersection with) the most voters (intervals). We only keep track of the last candidate ($lastc$) and the next candidate chosen ($nextc$), the number of candidates selected so far ($numc$), the number of voters hit so far ($numv$), and a counter to iterate through the voters ($i$). We also need to make sure that the maximum output value corresponds to a set of $k$ candidates that hits the most voters ($optv$). Our algorithm is presented in Algorithm 1.
Algorithm 1 SP-CC
@@
2:  append $01^{optv\cdot m(m+1)}$ 0 to output
@@
5:     guess $nextc$ such that $lastc<nextc\leq m$
@@
9:     append $01^{lastc}$ 0  to output
@@
Note that every interval that is hit is counted exactly once, namely the first time we choose a candidate that is in that interval. Line 7 ensures that later iterations do not count this interval again. Also note that the total length of the output produced in Line 9 is clearly upper bounded by $m(m+1)$, and so it follows that larger values of $optv$ will always give larger output values. Thus, the maximum output on an accepting path will correspond to a solution of the SP-CC problem. ∎
@@
For $1\leq i\leq m$, let $[s_{i},t_{i}]$ be the interval corresponding to the $i$th candidate in nondecreasing order of $s_{i}$, i.e., $s_{i}$ and $t_{i}$ are in $\{1,\ldots,n\}$ and the $i$th candidate (in an order where the candidates are ordered in nondecreasing order of $s_{i}$) is approved by the voters from the $s_{i}$th voter to the $t_{i}$th voter on our SC axis $L$. Note that we can’t store all these values, but we can recompute each value in logarithmic space whenever it is needed.
We need to compute a set of $k$ intervals (candidates) that cover the most voters, i.e., whose union is the largest. Note that we can’t simply guess $k$ intervals and compute the size of the union, since we do not have the space to store $k$ intervals. But we can limit what we need to store when going through the intervals in nondecreasing order of start ($s_{i}$). We only keep track of the last interval ($lastc$) and the next interval chosen ($nextc$), the number of intervals (candidates) selected so far ($numc$), the number of voters covered so far ($numv$), and the last voter covered ($lastv$). We also need to make sure that the largest output corresponds to a set of $k$ intervals that covers the most voters ($optv$). Our algorithm is presented in Algorithm 2.
Algorithm 2 SC-CC
@@
2:  append $01^{optv\cdot m(m+1)}$ 0 to output
@@
5:     guess $nextc$ such that $lastc<nextc\leq m$
@@
7:     $numv$ += $\left\|[lastv+1,n]\cap[s_{nextc},t_{nextc}]\right\|$
8:     append $01^{nextc}$ 0 to output
@@
Note that the total length of the output produced in Line 8 is clearly upper bounded by $m(m+1)$, and so it follows that larger values of $optv$ will always give larger output values. Thus, the maximum output on an accepting path will correspond to a solution of the SC-CC problem. ∎

### v3

@@ Abstract
Approval-Based Committee (ABC) rules are an important tool for choosing a fair set of candidates when given the preferences of a collection of voters. Though finding a winning committee for many ABC rules is NP-hard, natural variations for these rules with polynomial-time algorithms exist. The recently introduced Method of Equal Shares, an important ABC rule with desirable properties, is also computable in polynomial time. However, when working with very large elections, polynomial time is not enough and parallelization may be necessary. We show that computing a winning committee using these polynomial-time ABC rules (including the Method of Equal Shares) is P-hard, thus showing they cannot be parallelized. In contrast, we show that finding a winning committee can be parallelized when the votes are single-peaked or single-crossing for the important ABC rule Chamberlin-Courant.
@@ 1 Introduction and Related Work
Finding a winning committee for many ABC rules is NP-hard (Aziz et al., 2015; Brill et al., 2024; Godziszewski et al., 2021; LeGrand et al., 2007; Procaccia et al., 2008; Skowron et al., 2016). However, natural variations of these NP-hard rules exist where a winning set of candidates can be found in polynomial time. In addition, the Method of Equal Shares, which has desirable properties—such as satisfying a notion of group fairness known as Extended Justified Representation, can be computed in polynomial time (Peters and Skowron, 2020).
@@
It is important that we can always easily compute the outcome of an election. Previous work on studying the complexity of winner determination has largely focused on whether the winner problem is in P or NP-hard, equating being in P with being easy to compute. However, elections can be very large (e.g., Boehmer et al. (2024) note that blockchains that use the Nominated Proof-of-Stake protocol (e.g., Polkadot) use ABC rules to elect “validators” to ensure the integrity of the blockchain). In such cases it is not enough to just have a polynomial-time algorithm. In this case parallelization may be necessary.
@@
The parallelizability of voting rules was previously considered for single-winner voting rules (Csar et al., 2017, 2018), but not for the case of multiwinner rules. In fact, the recent textbook on ABC rules by Lackner and Skowron (2023) explicitly includes as open question “Q20” the task of determining which polynomial-time ABC rules are inherently sequential, i.e., do not have efficient parallel algorithms. As pointed out by Lackner and Skowron (2023), Approval Voting (AV) is clearly parallelizable (and the analogous result for Satisfaction Approval Voting (SAV) follows from the same argument). We show that computing the winning committee for all other studied polynomial-time ABC rules is inherently sequential.
@@
To show that a problem is inherently sequential we use the notion of P-hardness, since it is generally assumed that P-hard problems are not parallelizable (Greenlaw et al., 1995). We mention that this approach was used previously in computational social choice to show results about single-winner elections (specifically that winner determination for resolute versions of STV and Ranked Pairs is P-complete (Csar et al., 2017, 2018, 2025)), problems related to the aggregation of CP-nets (Lukasiewicz and Malizia, 2022), and for finding the essential set (Brandt and Fischer, 2008). We also mention a line of research that studies the parallelizability of problems in fair allocation Garg and Psomas (2023); Zheng and Garg (2019).
@@
When studying the complexity of a voting problem it is natural to consider settings where the votes of the electorate have structure. Two important domain restrictions are single-peaked and single-crossing preferences (Black, 1948; Mirrlees, 1971), which each model settings where the votes of the electorate are based on a single dimension (e.g., liberal vs. conservative). Both of these models have natural counterparts for approval voting (Elkind and Lackner, 2015; Faliszewski et al., 2011). We show that for single-peaked and for single-crossing approval votes finding a winning committee using the important Chamberlin-Courant ABC rule can be parallelized. We note here that the only previously-known parallelizability results for voting rules that we are aware of (in addition to AV and SAV mentioned previously (Lackner and Skowron, 2023)) are the one for Schulze Voting (Csar et al., 2018) as well as computing the Copeland, Smith, and Schwartz choice sets (Brandt et al., 2009).
@@
Inherently sequential ABC rules. We provide P-hardness reductions to prove the following theorem which encompasses all studied polynomial-time ABC rules from Lackner and Skowron (2023) except AV and SAV which are easily seen to be parallelizable (Lackner and Skowron, 2023); in fact they are computable in logarithmic space (see Section A of the appendix for completeness).
@@
Computing a winning committee is inherently sequential (P-hard) for seq-CC, seq-PAV, rev-seq-CC, rev-seq-PAV, seq-Phragmén, Greedy Monroe, and the Method of Equal Shares.
@@
Instead of providing parallel algorithms, we show membership in the relatively unknown complexity class OptL from Álvarez and Jenner (1993) (Definition  16). This allows for stronger results (since OptL is thought to be a strict subset of the class of parallelizable problems) and simpler proofs (since describing parallel algorithms can be an arduous task, e.g., one has to decide on a model and describe how each machine aggregates information, etc.).
@@ 2.1 Approval-Based Committee (ABC) Rules
Given candidates $C$, collection of votes $V$, and committee size $k$, output a committee $W$ of size $k$ such that it maximizes the number of voters that have at least one approved candidate in the committee.
@@ 2.2 Computational Complexity
Most of our results concern showing problems to be inherently sequential. It is widely believed that problems that are P-hard are inherently sequential (Greenlaw et al., 1995). Note that when showing a problem is hard for a given complexity class we must be careful to use a reduction with less computational power than the class we are showing hardness for. As is standard, we use logspace reductions for our P-hardness results, meaning that the reduction is computable in logarithmic space. [^1] Though logarithmic space may seem quite restrictive, standard NP-hardness reductions are computable in logarithmic space (see, e.g., Garey and Johnson (1979)).
@@
[…]
@@ 3 Inherently Sequential ABC Rules
Recall that finding the winning committee [^2] Note that the polynomial-time ABC rules from Lackner and Skowron (2023) explicitly include lexicographic tie-breaking, and so we can talk about the winning committee. However, we will also show that our hardness results go through when tie-breaking is never invoked, which shows that the hardness is not caused by tie-breaking. is a function problem. However, it gives more insight to prove the hardness of a related decision problem. Thus, we consider the decision problem that asks about the containment of a single candidate. First, we show that this formulation is correct by showing the equivalence of the decision and function problem with respect to parallelizability.
@@ 3.1 Chamberlin-Courant
While computing the winning committee for CC is NP-hard (Procaccia et al., 2008) in the general case, simple greedy approaches have been introduced to mimic CC in polynomial time. One such approach is the sequential Chamberlin-Courant rule, denoted seq-CC, where we start with an empty committee and iteratively add the candidate that increases the number of currently hit votes the most (see Lackner and Skowron (2023)).
@@
We will show that computing the winning committee using seq-CC is P-hard. Specifically, we show that deciding whether a given candidate is picked for the seq-CC winning committee is P-hard. To do this, we give a logspace reduction from a closely related P-complete problem, OVR, defined below. We note that as with NP-hardness reductions, half the battle is finding the right problem to reduce from, and we were able to find a problem almost equivalent to seq-CC. We hope that the straightforward nature of our reduction will ease the reader into P-hardness reductions.
@@
Figure 1: Column (a) shows the graph input to OVR at the top and the corresponding seq-$C$ election at the bottom. Columns (b) and (c) show the state of OVR and seq-$C$ after one and two rounds, respectively. Candidates are underlined when they are put in the winning committee and votes are darkened when they are hit. Observe how the removal of the highest-degree vertex in OVR corresponds to its inclusion in the winning committee in seq-$C$.
@@
As mentioned previously, the polynomial-time ABC rules from Lackner and Skowron (2023) explicitly include lexicographic tie-breaking, and so Theorem 7 proves that seq-CC in the definition of (Lackner and Skowron, 2023) is inherently sequential (P-hard). However, one might worry that the tie-breaking is causing the hardness. We now show that that is not the case.
Theorem 8.
Computing the winning committee for seq-CC is inherently sequential even when tie-breaking is never invoked.
Proof.
Suppose we are given an instance of seq-CC with candidates $C=\{c_{1},c_{2},\ldots,c_{m}\}$ in lexicographic order, a collection of votes $V$, and committee size $k$. We construct a new collection of votes, $V^{\prime}$, for which the same candidates as for $V$ will be selected, but for which tie-breaking is never invoked. Let $V^{\prime}$ be a collection of votes with $m$ copies of each vote in $V$. In addition, we add padding votes: for candidate $c_{i}$, we add mim-i votes that approve of only $c_{i}$. Note that the padding votes matter only in case of ties during the execution of seq-CC on the original election and so the elections select the same committee but the execution of seq-CC on the modified election never has ties. ∎
The rest of our hardness proofs reduce from a problem from which a straightforward reduction also introduces many ties. However, to avoid concerns about whether the hardness is linked to tie-breaking, we also provide proofs that ensure that tie-breaking is never invoked.
@@ 3.2 Method of Equal Shares
The Method of Equal Shares (MES) is a recently introduced rule (Peters and Skowron, 2020) that allots voters budgets and allows them to elect candidates using shares of their budget. A generalized version of this rule (Peters et al., 2021), has also been used in the real world for participatory budgeting (a setting that generalizes multiwinner elections to voting over issues to be funded) (see Peters and Skowron (2024)).
@@
[…]
Note that voters may run out of the budget necessary to elect another candidate before $k$ rounds. In this case, the method terminates early and returns the winning committee constructed so far. However, technically this does not meet our definition of an ABC rule, since we must return a committee of size $k$. To rectify this, another ABC rule is called to select the remaining candidates—this is referred to as the second phase of MES. Lackner and Skowron (2023) suggest using seq-Phragmén (see Sections C.2 and C.4 of the appendix) since it preserves some desirable properties.
We will show that computing the committee found by (the first phase of) MES is P-hard. This shows that the hardness of MES is not caused by the hardness of the rule called afterward (such as seq-Phragmén). As in the previous section, we focus on the decision problem where we ask about the inclusion of a specified candidate. We reduce from a problem known as LFMIS:
Problem 11 (Lexicographically First Maximal Independent Set (LFMIS) (Cook, 1985)).
@@
Computing the winning committee for (the first phase of) MES is inherently sequential.
@@
Suppose we are given an instance $(G,v)$ of LFMIS, where $G$ is a 3-regular graph with $m$ vertices labeled $1$ to $m$. We will construct an election with candidates $C$, collection of votes $V$, and committee size $k$, and specify a candidate $c\in C$ such that the winning committee selected by MES includes $c$ if and only if $v$ is included in the LFMIS of $G$.
@@
We note that the above reduction may return a winning committee with fewer than $k$ candidates. Recall that in this case another ABC rule is called to elect the remaining candidates. For the sake of completeness, we show in Section C.4 of our appendix that MES is also P-hard when seq-Phragmén is used, as suggested by Lackner and Skowron (2023), to finish electing a total of $k$ candidates. Particularly, we reduce from LFMIS and show that the specified vertex $v$ is picked by MES if and only if $v$ belongs to the LFMIS.
As in the case of seq-CC we can show that (the first phase of) MES is P-hard for the case where tie-breaking is never invoked. Informally, we pad our reduction from LFMIS with enough votes so that none of the candidates are tied initially, and through the execution of MES at no point will candidates become tied. See Section C.5 of the appendix for details.
Theorem 13.
Computing the winning committee for (the first phase of) MES is inherently sequential even when tie-breaking is never invoked.
@@ 4 Domain Restrictions
Figure 3: Column (a) shows an election. Column (b) shows how the votes are single-peaked with respect to the axis cbadc\ b\ a\ d. Column (c) shows how the votes are single-crossing with respect to the axis 1 2 3 41\ 2\ 3\ 4. The intervals (candidates) in Column (c) are ordered in nondecreasing order with respect to starting point, as will be done in the construction of Theorem 18.
@@
As pointed out by Faliszewski et al. (2011) (for single-peaked) and Elkind and Lackner (2015) (for single-crossing), the problem of determining whether a collection of votes (in our setting of approval ballots) is single-peaked or single-crossing, and computing an axis witnessing this is in essence the problem of determining whether a 0-1 matrix has the consecutive ones property and computing a permutation that witnesses that. These problems have long been known to be computable in polynomial time (Booth and Lueker, 1976; Fulkerson and Gross, 1965).
@@
We provide a visualization in Figure 4. It is known that every OptL function is parallelizable (Álvarez and Jenner, 1993, Theorem 4.2). We will show that SP-CC and SC-CC are in OptL (Theorems 17 and 18). Since OptL is a subset of the class of parallelizable problems, this immediately implies Theorem 15. As an aside, a benefit of this approach is that we can show parallelizability without needing to directly describe a parallel algorithm, which can be quite involved (e.g., we would first need to decide on a model, describe how each machine aggregates information, we would need to show polylogarithmic running time, etc).
@@
For $1\leq i\leq n$, let $s_{i},t_{i}]$ be the interval corresponding to the $i$th voter, i.e., $s_{i}$ and $t_{i}$ are in $\{1,\ldots,m\}$ and the $i$th voter approves the candidates from the $s_{i}$th candidate to the $t_{i}$th candidate on our SP axis $L$. Note that we can’t store all these values, but we can recompute each value in logarithmic space whenever it is needed.
We now need to compute a set of $k$ candidates that hit (have nonempty intersection with) the most voters (intervals). We only keep track of the last candidate (lastclastc) and the next candidate chosen (nextcnextc), the number of candidates selected so far (numcnumc), the number of voters hit so far (numvnumv), and a counter to iterate through the voters ($i$). We also need to make sure that the maximum output value corresponds to a set of $k$ candidates that hits the most voters (optvoptv). Our algorithm is presented in Algorithm 1.
Algorithm 1 SP-$C$
@@
2:append $1^{optv\cdot m(m+1)}0$ to output
@@
5:  guess nextcnextc such that $lastc<nextc\leq m$
@@
9:  append $1^{lastc}0$  to output
@@
Note that every interval that is hit is counted exactly once, namely the first time we choose a candidate that is in that interval. Line 7 ensures that later iterations do not count this interval again. Also note that the total length of the output produced in Line 9 is clearly upper bounded by $m(m+1)$, and so it follows that larger values of optvoptv will always give larger output values. Thus, the maximum output on an accepting path will correspond to a solution of the SP-CC problem. ∎
@@
For $1\leq i\leq m$, let $s_{i},t_{i}]$ be the interval corresponding to the $i$th candidate in nondecreasing order of $s_{i}$, i.e., $s_{i}$ and $t_{i}$ are in $\{1,\ldots,n\}$ and the $i$th candidate (in an order where the candidates are ordered in nondecreasing order of $s_{i}$) is approved by the voters from the $s_{i}$th voter to the $t_{i}$th voter on our SC axis $L$. Note that we can’t store all these values, but we can recompute each value in logarithmic space whenever it is needed.
We need to compute a set of $k$ intervals (candidates) that cover the most voters, i.e., whose union is the largest. Note that we can’t simply guess $k$ intervals and compute the size of the union, since we do not have the space to store $k$ intervals. But we can limit what we need to store when going through the intervals in nondecreasing order of start ($s_{i}$). We only keep track of the last interval (lastclastc) and the next interval chosen (nextcnextc), the number of intervals (candidates) selected so far (numcnumc), the number of voters covered so far (numvnumv), and the last voter covered (lastvlastv). We also need to make sure that the largest output corresponds to a set of $k$ intervals that covers the most voters (optvoptv). Our algorithm is presented in Algorithm 2.
Algorithm 2 SC-$C$
@@
2:append $1^{optv\cdot m(m+1)}0$ to output
@@
5:  guess nextcnextc such that $lastc<nextc\leq m$
@@
7:  numvnumv += $\left\|[lastv+1,n]\cap[s_{nextc},t_{nextc}]\right\|$
8:  append $1^{nextc}0$ to output
@@
Note that the total length of the output produced in Line 8 is clearly upper bounded by $m(m+1)$, and so it follows that larger values of optvoptv will always give larger output values. Thus, the maximum output on an accepting path will correspond to a solution of the SC-CC problem. ∎